the value of x in the equation 6x-5(2+2x)=x-10 is
Answer by Theo(13342)

6x-5*(2+2x)=x-10
simplify to get:
6x-10-10x=x-10
combine like terms to get:
-4x-10=x-10
add 10 to both sides of the equation and subtract x from both sides of the equation to get:
-5x=0
solve for x to get:
x=0.
the two equations are equal when x = 0.
when x = 0, .....
6x-5*(2+2x) becomes -5*2 which becomes -10
x-10 becomes -10
if you let y1 = 6x-5*(2+2x) and you let y2 = x-10, y1 will be equal to y2 when x = 0
if you graph the equations, you will find that the two lines formed intersect at the coordinate point of (x,y) = (0,-10).
here's the graph: